You drop a rock from a 6th story window. At time t it passes the 5th story window. What window will it be passing at time 2t? Assume the windows are equally spaced.

A. 2nd story

B. 5th story

C. 4th story

D. 3rd story
Answer #1

Solution:

First condiction, Distace nedd to be covered for 1 floor

d = 1/2gt2

2nd condiction,

d' = 1/2 g(2t)2

d' = 1/2 g *4t2

d' = 4*[1/2gt2] = 4d

Ans : A. 2nd story.

I hope you understood the problem, If yes rate me!! or else comment for a better solution